OpenStudy (anonymous):

Which rational number is one-half of the way between 2 and 6.5: 2.25, 4.25, 4.5, or 4.75? And please explain.

OpenStudy (barrycarter):

To find the average, add and divide by 2, so (2+6.5)/2 = 4.25

OpenStudy (anonymous):

The way between 2 and 6.5 is 4.5. Half of that way is 2.25.

OpenStudy (anonymous):

We have 2 numbers: 2 and 6.5 One-half means the answer divides that into 2 parts. So, you need to find the distance between 6.5 and 2, so 6.5-2 = 4.5 Half of 4.5 is 4.5/2 = 2.25 Got it? :)